Problem 14-04
Stock Repurchase

A firm has 10 million shares outstanding with a market price of $35 per share. The firm has $35 million in extra cash (short-term investments) that it plans to use in a stock repurchase; the firm has no other financial investments or any debt. What is the firm's value of operations after the repurchase? Enter your answer in millions. For example, an answer of $1.2 million should be entered as 1.2, not 1,200,000. Round your answer to two decimal places.

$ million

How many shares will remain after the repurchase? Round your answer to the nearest whole number.

shares
Answer #1

Firm value after repurchase = price*shares-cash = 10*35-35 =315m

Share after repurchase = current shares-cash/price = 10-35/35=9m